[figure: 9.7 cm]

COOPERATIVE JAHN-TELLER EFFECT IN LOW-SYMMETRY

RARE-EARTH SYSTEMS

Henryk Szymczak Institute of Physics, Polish Academy of Sciences, Warsaw, Poland

The cooperative Jahn-Teller effect (CJTE) has many fascinating aspects related mainly to structural phase transitions. In the CJTE-systems the driving force for the phase transition is the coupling between the lattice deformations and the electronic degrees of freedom. This coupling can be considered as an effective quadrupole-quadrupole interaction mediated by phonons. The Jahn-Teller effect is generally quite small in rare-earth compounds presumably because of the smallness of the electron-phonon coupling. Recently we have shown that in some low-symmetry rare-earth systems this interaction is relatively strong and CJTE can be observed by means of a variety of experimental methods. In this review the results of structural, magnetic, optical, EPR and specific heat studies of CJTE in rare-earth double tungstates and double molybdates are presented. These materials crystallize in low-symmetry structures and many of them show complicated structural and magnetic phase transitions driven by CJTE. It is shown that molecular-field approximation gives satisfactory description of both the static and dynamic behavior of these systems.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : PLENARY LECTURE

Corresponding author : H. Szymczak

Address for correspondence : Instytut Fizyki PAN Al. Lotników 32/46 02–668 Warszawa

Email address : szymh@ifpan.edu.pl